OpenStudy (anonymous):

Multiply and simplify: b2 • b4 • b • b5

OpenStudy (kc_kennylau):

I suppose you mean \(b^2\times b^4\times b\times b^5\)?

OpenStudy (cggurumanjunath):

just add the powers

OpenStudy (cggurumanjunath):

b^ (2+4+1+5)